# Elucidating the Design Space of Arbitrary-Noise-Based Diffusion Models

### v1

@@ Abstract
EDM elucidates the unified design space of diffusion models, yet its fixed noise patterns restricted to pure Gaussian noise, limit advancements in image restoration. Our study indicates that forcibly injecting Gaussian noise corrupts the degraded images, overextends the image transformation distance, and increases restoration complexity. To address this problem, our proposed EDA Elucidates the Design space of Arbitrary-noise-based diffusion models. Theoretically, EDA expands the freedom of noise pattern while preserving the original module flexibility of EDM, with rigorous proof that increased noise complexity incurs no additional computational overhead during restoration. EDA is validated on three typical tasks: MRI bias field correction (global smooth noise), CT metal artifact reduction (global sharp noise), and natural image shadow removal (local boundary-aware noise). With only 5 sampling steps, EDA outperforms most task-specific methods and achieves state-of-the-art performance in bias field correction and shadow removal.
Figure 1: Our EDA supports arbitrary noise patterns diffusion ($\text{N}_{\text{Img}}$) and enables the reverse process to be initiated directly from the known degraded image. This avoids the extra Gaussian noise corruption ($\text{N}_{\text{Gaus}}$; top row) of EDM-based methods, and shortens the image transformation distance, while reducing task complexity and achieving high-quality results with fewer sampling steps.
@@ 1 Introduction
Existing diffusion-based generative models have been proven to incorporate into a unified design space known as EDM [8]. However, EDM is limited to Gaussian noise diffusion, and does not encompass methods with other patterns of noise diffusion process [13, 1]. While EDM unifies diverse theoretical frameworks [21, 6, 22, 23] into a shared design space with flexible structural parameters, such as noise schedules and training objectives [8], its reliance on pure Gaussian noise prevents integration of methods like Flow Matching [13] and Cold Diffusion [1] that support diverse diffusion processes between arbitrary images as shown in Fig. 2(a). For example, transforming images $A$ to $B$ requires extending the diffused noise patterns beyond Gaussian noise to residual domains ($B-A$). Consequently, increasing the flexibility of noise patterns and constructing a more inclusive design space remains an open challenge.
@@
(b) Comparison of transformation distance for diffusion.
Figure 2: (a) EDA extends EDM by enabling any diffused noise patterns while retaining the flexibility of structural parameters, such as noise schedules and training objectives. (b) EDA enables arbitrary noise diffusion avoiding extra Gaussian noise, and reducing the image transformation distance during the reverse process.
High-flexibility diffused noise patterns provide greater advantages for image restoration compared to generation. When restricted to Gaussian noise, vanilla diffusion-based restoration methods [14, 16, 17, 9] initiate the reverse process from noise-corrupted degraded images or pure noise maps, progressively restoring them toward the target domain. However, this forced noise injection introduces two key limitations: (1) degraded images lose task-specific information due to added noise corruption, (2) the imposed noise artificially increases the image transformation distance and restoration complexity as shown in Fig. 2(b). This distance can be reduced by customizing the diffused noise patterns and initiating the reverse process directly from the known degraded image in the task, thereby simplifying the restoration. In contrast, generation tasks lack the degraded image to exploit. Consequently, high-flexibility diffused noise patterns are critical for restoration task but remain unaddressed in the unified EDM framework [8].
This paper elucidates the design space of diffusion models with arbitrary noise patterns, termed EDA. EDA characterizes the diffusion process through multivariate Gaussian distributions, and our study theoretically demonstrates its capacity to diffuse arbitrary noise patterns. The framework models diffused noise via multivariate Gaussian distributions and derives a corresponding stochastic differential equation (SDE) driven by multiple independent Wiener processes. Deterministic sampling rules are derived by solving the probability flow ordinary differential equation (PFODE). Our study indicates that EDA incurs no additional computational cost when generalizing from simple Gaussian noise to complex patterns. EDA is evaluated on three restoration tasks: MRI bias field correction (global smooth noise), CT metal artifact reduction (global sharp noise), and natural image shadow removal (local boundary-aware noise). With fewer than $5$ sampling steps, EDA performs competitively with $100$-step EDM-based methods, outperforms task-specific methods, and establishes state-of-the-art results in bias field correction and shadow removal.
@@ 2.1 EDM can Only Diffuse Gaussian Noise
$\mathrm{d}\bm{x}=f(t)\bm{x}\,\mathrm{d}t+g(t)\,\mathrm{d}\omega_{t},$ (1)
@@
While EDM offers theoretical unification, its single diffused noise pattern introduces critical limitations. (1) The closed-form perturbation kernel in Eq. 2 limits diffusion to pure Gaussian noise characterized by pixel-wise independence and homogeneity, precluding support for arbitrary noise patterns. (2) In restoration tasks where the initial state $x_{T}$ is the degraded image $P_{LQ}$, however EDM-based methods [16, 9, 17, 14] introduces an artificial restoration gap as its initial state must include Gaussian noise $P_{LQ}+N_{Gaus}$. Our EDA addresses this by increasing the flexibility of diffused noise patterns while retaining the flexibility of structural parameters.
@@ 2.2 Diffusion-based Methods for Any Noise Diffusion
Recent advances in diffusion models have extended diffused noise from pure Gaussian noise to arbitrary noise patterns, but EDM as a unified design space cannot contain them. Flow Matching [13] enables transformations between arbitrary distributions by constructing continuous probability paths, directly modeling deterministic transport from an initial noise distribution $p_{1}(x)$ to a target data distribution $p_{0}(x)$ through conditional paths $\pi_{t}(x|x_{0})$. Its key innovation lies in designing vector fields via continuous normalizing flows (CNFs), allowing samples $x_{0}\sim p_{0}$ to evolve along these paths toward $x_{1}$. In contrast, Cold Diffusion [1] replaces Gaussian noise with deterministic degradation operators $D(\cdot)$, and trains networks to invert $D(\cdot)$ through restoration operators $R(\cdot)$.
These works demonstrate that the strength of diffusion models lies in their progressive transformation framework, not specific noise types. The probability paths of Flow Matching and deterministic operators of Cold Diffusion illustrate diverse diffusion processes. Following previous work, this paper denotes these diffusion mediums as generalized noise. Our EDA elucidates a design space that encompasses diffusion models with arbitrary noise patterns, positioning EDM-based methods as specific cases confined to pure Gaussian noise.
@@ 3.1 Generalized Forward Process
$\bm{x}_{t}=s(t)\bm{x}_{0}+s(t)\sigma(t)\sum_{m=1}^{M}\frac{\epsilon_{m}+\eta}{\eta+1}h_{m,\bm{x}_{0}},$ (9)
@@
$P(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}\left(\bm{x}_{t};\,s(t)\bm{x}_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}0},\,\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\bm{\Sigma}_{\bm{x}_{0}}\right)$ (10)
@@
$f(t)=\frac{s^{\prime}(t)}{s(t)},\quad g(t)=\frac{s(t)}{\eta+1}\sqrt{\frac{\mathrm{d}\sigma^{2}(t)}{\mathrm{d}t}},\quad\phi_{\bm{x}_{0}}(t)=\frac{\eta s(t)\sigma^{\prime}(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}$ (12)
@@ 3.2 Deterministic Sampling via PFODE
$\nabla{\bm{x}}\log P(\bm{x}|\bm{x}_{0})=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\bm{\Sigma}_{\bm{x}_{0}}^{-1}\left[s(t)\bm{x}_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-\bm{x}\right]$ (14)
@@
Algorithm 1 The EDA training process with arbitrary diffused noise pattern $N$.
1:$P_{data}$, $H$, $\eta$, $\sigma(t)$, $s(t)$, time training strategies $P_{time}$, training step $N$,
@@
8:     Sample diffusion data $\bm{x}_{t}\sim\mathcal{N}\left(s(t)\bm{x}_{0}+\frac{\eta s(t)\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}0},\,\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\bm{\Sigma}_{\bm{x}_{0}}\right)$
9:     Compute loss function:
$\mathcal{L}=\lVert D_{\theta}(\bm{x};\sigma)-\bm{x_{0}}\rVert^{2}$
10:     Update network parameters with $\nabla_{\theta}\mathcal{L}$
11:end for
12:return $D_{\theta}(\cdot)$triangleright Return trained denoiser
@@
1:$D_{\theta}(x;\sigma)$, $\sigma(t)$, $s(t)$, time steps $t_{i\in\{0,\ldots,N\}}$
@@
8:return $x_{0}$triangleright Return noise-free sample
Proposition 1.
EDA Supports Diffusion and Removal of Arbitrary Noise.
  1. Case 1
  2. Case 2
    Sample-Dependent Basis Functions (General Case). This case applies universally. When noise patterns cannot be decomposed into a fixed basis set, $H_{\bm{x}_{0}}$ becomes sample-dependent. Take a simple example, mapping an image $\bm{A}$ to its degraded counterpart $\bm{B}$, where the basis set reduces to $H_{\bm{A}}=[\bm{B}-\bm{A}]$. Unlike Case 1, Case 2 uses the approximately optimal output of the denoiser in the deterministic sampling process (see Appendix C). Applications include metal artifact reduction (MAR) in CT and shadow removal in natural images (Appendices E.2, E.3).
  3. Case 3
@@
Proof. EDA aligns with EDM in deterministic sampling and denoiser training objective. While the diffusion process in EDA incorporates an arbitrary covariance matrix $\Sigma_{\bm{x}_{0}}$, contrasting with EDM’s diagonal covariance $\sigma^{2}(t)\mathbf{I}$, this introduces additional terms in the $P(\bm{x}_{t}|\bm{x}_{0})$ (Eq. 2 vs. 10), the SDE (Eq. 1 vs. 11), the PFODE (Eq. 3 vs. 13), and the score functions (Eq. 4 vs. 14). Remarkably, these additional terms can be analytically simplified and eliminated (Appendix C), resulting in our deterministic sampling formulation Eq. 15) that aligns with EDM (Eq. 7). Furthermore, the denoiser is trained using the same objective, as shown in Eq. 5 and 16.
@@
Proof. EDM [8] is a strict subset of EDA under specific parameter configurations as shown in Fig. 2(a). When the parameters are set to: (1) the mediator $\eta=0$; (2) the basis set $H=[E_{1,1},...,E_{H,W}]$, where each $E_{i,j}$ is a matrix with a value of $1$ at position $(i,j)$ and 00 elsewhere. The diffused noise $N$ in EDA Eq. 8 reduces to pixel-wise independent Gaussian noise:
@@ 4.1 Tasks and Datasets
Image- domain Refusion 33.66/0.9694 33.82/0.9748 39.36/0.9851 43.58/0.9858 40.31/0.9812 38.15/0.9793
@@ 4.2 Experimental Setup
Our EDA was developed in PyTorch and trained on a single NVIDIA GeForce RTX 3090 GPU. The parameters are set as, $s(t)=1$, $\sigma=\sqrt{1-\bar{\alpha_{t}}}$, where ${\bar{\alpha_{t}}}$ is hyperparameter as in the DDPM [6]. The total step during training is $T=100$. The training time strategies are denoted as $P_{time}\sim U(0,T)$. For BFC, the mediator $\eta=0$, the basis set $H$ was set as Case 1 in Prop. 1. The input MRIs were $256\times 256$ and trained for $500$ epochs with a batch size of $1$. For MAR and SR, the $\eta=10$ and the basis set $H$ followed Case 2 in Prop. 1. The input CTs were $416\times 416$ and trained $1000$ epochs with a batch size of $1$. The Shadow Images were $256\times 256\times 3$, trained $2000$ epochs with a batch size of $4$. The specific experimental setup is shown in the Appendix E.
Our EDA was compared with task-specific methods. The BFC methods included N4 [24], MICO [10], LapGM [26], ABCNet [2]. The MAR included LI [7], CNNMAR [33], DSCMAR [31], DuDoNet [12], InDuDoNet+ [28], DICDNet [27], which are all uses of dual domains (sinusoidal and image domain). The SR included G2R [15], DHAN [3], Fu et al. [4], Zhu et al. [35], BMNet [34], ShadowFormer [5].
Furthermore, the experiments included Refusion [17], a representative Gaussian diffusion-based restoration method, and an optimized version of IRSDE [16]. Refusion secured second place overall (first in perceptual metrics) in the CVPR NTIRE Shadow Removal Challenge. All results were either quoted from original publications or reproduced using their official implementations, with Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) as the primary evaluation metrics.

### v2

@@ Abstract
Although EDM aims to unify the design space of diffusion models, its reliance on fixed Gaussian noise prevents it from explaining emerging flow-based methods that diffuse arbitrary noise. Moreover, our study reveals that EDM’s forcible injection of Gaussian noise has adverse effects on image restoration task, as it corrupts the degraded images, overextends the restoration distance, and increases the task’s complexity. To interpret diverse methods for handling distinct noise patterns within a unified theoretical framework and to minimize the restoration distance, we propose EDA, which Elucidates the Design space of Arbitrary-noise diffusion models. Theoretically, EDA expands noise pattern flexibility while preserving EDM’s modularity, with rigorous proof that increased noise complexity introduces no additional computational overhead during restoration. EDA is validated on three representative medical image denoising and natural image restoration tasks: MRI bias field correction (global smooth noise), CT metal artifact removal (global sharp noise) and natural image shadow removal (local boundary-aware noise). With only 5 sampling steps, competitive results against specialized methods across medical and natural tasks demonstrate EDA’s strong generalization capability for image restoration. Code is available at: https://github.com/PerceptionComputingLab/EDA.
@@ 1 Introduction
Figure 1: Our EDA supports arbitrary noise patterns diffusion ($\text{N}_{\text{Img}}$) and enables the reverse process to be initiated directly from the known degraded image. This avoids the extra Gaussian noise corruption ($\text{N}_{\text{Gaus}}$; top row) of EDM-based methods, and shortens the image restoration distance, while reducing task complexity and achieving high-quality results with fewer sampling steps.
Currently, there is no unified design space for diffusion models that accommodates both flexible noise patterns and high-performance SDE-based diffusion, which hinders the theoretical development of more generalized diffusion models. EDM [11] unifies most diffusion models [28, 9, 29, 30], but its restriction to diffusing only Gaussian noise (as shown in Fig. 2(a)) prevents it from encompassing more flexible diffusion methods [1, 7]. Alternatively, Flow Matching [17] offers a ODE diffusion framework, breaking the constraints of Gaussian noise diffusion. This allows more flexible diffusion processes and enables more efficient sampling. Nevertheless, SDE-based approaches have superior diversity and quality of results as shown in numerous studies [28, 30, 11] compared to ODE diffusion. Consequently, establishing a unified SDE-based design space that enhances the flexibility of noise patterns is important, which would empower researchers to articulate the theoretical distinctions between disparate diffusion methods from a consistent viewpoint, thereby laying a robust foundation for future diffusion model development. Addressing this open challenge, the EDA diffusion framework proposed in this paper offers a viable solution.
@@
(b) Comparison of restoration distance for diffusion.
Figure 2: (a) EDA extends EDM by enabling any diffused noise patterns while retaining the flexibility of structural parameters, such as noise schedules and training objectives. (b) EDA enables arbitrary noise diffusion avoiding extra Gaussian noise, and reducing the image restoration distance during the reverse process.
High-flexibility diffused noise patterns provide greater advantages for image restoration compared to generation. When restricted to Gaussian noise, vanilla diffusion-based restoration methods [19, 22, 21, 12] initiate the reverse process from noise-corrupted degraded images or pure noise maps, progressively restoring them toward the target domain. However, this forced noise injection introduces two key limitations: (1) degraded images lose task-specific information due to added noise corruption, (2) the imposed noise artificially increases the image restoration distance and restoration complexity as shown in Fig. 2(b). This distance can be reduced by customizing the diffused noise patterns and initiating the reverse process directly from the known degraded image in the task, thereby simplifying the restoration. Consequently, the restrictive reliance of the EDM framework on purely Gaussian diffusion noise fundamentally undermines its efficiency and practical utility for high-fidelity image restoration tasks.
This paper Elucidates the Design space of diffusion models with Arbitrary noise patterns, termed EDA. EDA interprets diffusion methods handling diverse noise patterns within a unified theoretical framework as shown in Fig. 1, simplifying and unifying diffusion-based restoration processes originally developed for Gaussian noise constraints. Specifically, EDA characterizes the diffusion process through multivariate Gaussian distributions, and our study theoretically demonstrates its capacity to diffuse arbitrary noise patterns. The framework models diffused noise via multivariate Gaussian distributions and derives a corresponding stochastic differential equation (SDE) driven by multiple independent Wiener processes. Deterministic sampling rules are derived by solving the probability flow ordinary differential equation (PFODE). Our study indicates that EDA incurs no additional computational cost when generalizing from simple Gaussian noise to complex patterns. EDA is evaluated on three restoration tasks, including two critical medical image denoising applications: MRI bias correction (global smooth noise) and CT metal artifact reduction (global sharp noise), along with natural image shadow removal (local boundary-aware noise), which is commonly used. Superior experimental results on both medical and natural images demonstrate EDA’s strong generalization ability for image restoration. With fewer than $5$ sampling steps, EDA performs competitively with $100$-step EDM-based methods, outperforms task-specific methods, and establishes state-of-the-art results in bias field correction and shadow removal.
@@ 2.1 EDM can Only Diffuse Gaussian Noise
$\mathrm{d}\bm{x}=f(t)\bm{x}\,\mathrm{d}t+g(t)\,\mathrm{d}\omega_{t}$ (1)
@@
While EDM offers theoretical unification, its single diffused noise pattern introduces critical limitations. (1) The closed-form perturbation kernel in Eq. 2 limits diffusion to pure Gaussian noise characterized by pixel-wise independence and homogeneity, precluding support for arbitrary noise patterns. (2) In restoration tasks where the initial state $x_{T}$ is the degraded image $P_{LQ}$, however EDM-based methods [22, 12, 21, 19, 18, 4, 13, 41, 5, 27] introduces an artificial restoration gap as its initial state must include Gaussian noise $P_{LQ}+N_{Gaus}$. Our EDA addresses this by increasing the flexibility of diffused noise patterns while retaining the flexibility of structural parameters.
@@ 2.2 Diffusion-based Methods for Any Noise Diffusion
Recent advances in diffusion models have extended diffused noise from pure Gaussian noise to arbitrary noise patterns, but this comes at the cost of theoretical simplification, forcing the generation process to degrade from SDE to deterministic ODEs, or even lacking a clear theoretical grounding, which fundamentally prevents the EDM design space from encompassing these new approaches. Flow Matching [17] enables transformations between arbitrary distributions by constructing continuous probability paths, directly modeling deterministic transport from an initial noise distribution $p_{1}(x)$ to a target data distribution $p_{0}(x)$. Its key innovation lies in designing vector fields via continuous normalizing flows (CNFs), allowing samples $x_{0}\sim p_{0}$ to evolve along these paths toward $x_{1}$. The core dynamics are defined via ODE:
$\frac{dx_{t}}{dt}=v_{t}(x_{t})$ (8)
where $x_{t}$ denotes the sample state at time $t\in[0,1]$, and $v_{t}(x_{t})$ is the time-dependent velocity field learned by the model to drive this evolution. MeanFlow [7] extends Flow Matching by characterizing the flow field using the notion of average velocity, in contrast to the instantaneous velocity modeled by its predecessor. This theoretical refinement is crucial for efficiency, enabling MeanFlow to achieve state-of-the-art performance in one-step generative modeling, making it a critical benchmark for highly efficient diffusion and flow methods. Cold Diffusion [1] sacrifices the theoretical grounding in stochastic processes to gain enhanced flexibility and generality across tasks. It does so by replacing Gaussian noise with deterministic degradation operators $D(\cdot)$ and training networks to invert them via restoration functions $R(\cdot)$.
These works collectively propose diverse diffusion processes, demonstrating that the strength of diffusion models lies in their progressive transformation framework, not specific noise types. Despite the high flexibility of ODE diffusions, they may exhibit reduced robustness due to the lack of stochasticity [28, 30, 11], so the development of a unified framework that retains the SDE design space is essential for leveraging its advantages in generating high-fidelity and diverse samples. Following previous work, this paper denotes these diffusion mediums as generalized noise. Our EDA elucidates a comprehensive design space that encompasses both SDE-based diffusion models and diverse, generalized diffusion processes, thereby avoiding the restrictive Gaussian noise assumption inherent in EDM.
@@ 3.1 Generalized Forward Process
$\bm{x}_{t}=s(t)\bm{x}_{0}+s(t)\sigma(t)\sum_{m=1}^{M}\frac{\epsilon_{m}+\eta}{\eta+1}h_{m,\bm{x}_{0}}$ (10)
@@
$P(\bm{x}_{t}\mid\bm{x}_{0})$ $=\mathcal{N}\bigl(\bm{x}_{t};\,\mu_{t},\,\sigma_{t}\bigr),$ (11)
$\mu_{t}$ $=s(t)\,\bm{x}_{0}+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},$
$\sigma_{t}$ $=\frac{s^{2}(t)\,\sigma^{2}(t)}{(\eta+1)^{2}}\,\bm{\Sigma}_{\bm{x}_{0}}$
@@
[…]
@@ 3.2 Deterministic Sampling via PFODE
$\nabla_{\bm{x}}\log P(\bm{x}\mid\bm{x}_{0})=\frac{(\eta+1)^{2}}{s^{2}(t)\sigma^{2}(t)}\,\bm{\Sigma}_{\bm{x}_{0}}^{-1}$ (15)
$\quad\cdot\Bigl(s(t)\,\bm{x}_{0}+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}}-\bm{x}\Bigr)$
@@
Algorithm 1 The EDA training process with arbitrary diffused noise pattern NN.
[…]
@@
8:  Sample diffusion data: $\bm{x}_{t}\sim$
9:
$\mathcal{N}\!\left(s(t)\bm{x}_{0}+\frac{\eta\,s(t)\,\sigma(t)}{\eta+1}\sum_{m=1}^{M}h_{m,\bm{x}_{0}},\;\frac{s^{2}(t)\sigma^{2}(t)}{(\eta+1)^{2}}\bm{\Sigma}_{\bm{x}_{0}}\right)$
10:  Compute loss function: $\mathcal{L}=\lVert D_{\theta}(\bm{x};\sigma)-\bm{x_{0}}\rVert^{2}$
11:  Update network parameters with $\nabla_{\theta}\mathcal{L}$
12:end for
13:return $D_{\theta}(\cdot)$ $\triangleright$ Return trained denoiser
Proposition 1.
EDA Supports Diffusion and Removal of Arbitrary Noise.
  1. 1.
  2. 2.
    Sample-Dependent Basis Functions (General Case). This case applies universally. When noise patterns cannot be decomposed into a fixed basis set, $H_{\bm{x}_{0}}$ becomes sample-dependent. Take a simple example, mapping an image $\bm{A}$ to its degraded counterpart $\bm{B}$, where the basis set reduces to $H_{\bm{A}}=[\bm{B}-\bm{A}]$. Unlike Case 1, Case 2 uses the approximately optimal output of the denoiser in the deterministic sampling process (see Appendix C). Applications include metal artifact reduction (MAR) in CT and shadow removal in natural images (Appendices E.3, E.4).
  3. 3.
@@
[…]
@@
8:return $x_{0}$ $\triangleright$ Return noise-free sample
@@
Proof. EDA aligns with EDM in deterministic sampling and denoiser training objective. While the diffusion process in EDA incorporates an arbitrary covariance matrix $\Sigma_{\bm{x}_{0}}$, contrasting with EDM’s diagonal covariance $\sigma^{2}(t)\mathbf{I}$, this introduces additional terms in the $P(\bm{x}_{t}|\bm{x}_{0})$ (Eq. 2 vs. 11), the SDE (Eq. 1 vs. 12), the PFODE (Eq. 3 vs. 14), and the score functions (Eq. 4 vs. 15). Remarkably, these additional terms can be analytically simplified and eliminated (Appendix C), resulting in our deterministic sampling formulation Eq. 16) that aligns with EDM (Eq. 7). Furthermore, the denoiser is trained using the same objective, as shown in Eq. 5 and 17.
@@
Proof. EDM [11] is a strict subset of EDA under specific parameter configurations as shown in Fig. 2(a). When the parameters are set to: (1) the mediator $\eta=0$; (2) the basis set $H=[E_{1,1},...,E_{H,W}]$, where each $E_{i,j}$ is a matrix with a value of $1$ at position $(i,j)$ and 0 elsewhere. The diffused noise $N$ in EDA Eq. 9 reduces to pixel-wise independent Gaussian noise:
@@ 4.1 Tasks and Datasets
Image- domain MeanFlow 28.47/0.8795 31.09/0.9253 34.92/0.9391 37.01/0.9454 37.83/0.9489 33.86/0.9276
Refusion 33.66/0.9694 33.82/0.9748 39.36/0.9851 43.58/0.9858 40.31/0.9812 38.15/0.9793
@@ 4.2 Experimental Setup
Our EDA was developed in PyTorch and trained on a single NVIDIA GeForce RTX 3090 GPU. The parameters are set as, $s(t)=1$, $\sigma=\sqrt{1-\bar{\alpha_{t}}}$, where $\bar{\alpha_{t}}$} is hyperparameter as in the DDPM [9]. The total step during training is $T=100$. The training time strategies are denoted as $P_{time}\sim U(0,T)$. Task-specific training settings (e.g., input resolutions, epochs, and batch sizes) are provided in the Supplementary Material E.
Our EDA was compared with task-specific methods. The BFC methods included N4 [31], MICO [14], LapGM [33], ABCNet [2]. The MAR included LI [10], CNNMAR [40], DSCMAR [38], DuDoNet [16], InDuDoNet+ [35], DICDNet [34], which are all uses of dual domains (sinusoidal and image domain). The SR included G2R [20], DHAN [3], Fu et al. [6], Zhu et al. [43], BMNet [42], ShadowFormer [8]. Furthermore, our experiments involved representative diffusion methods, including Refusion [21], MeanFlow [7], and an optimized version of IRSDE [22]. Refusion is a representative Gaussian diffusion-based restoration method which secured second place overall (first in perceptual metrics) in the CVPR NTIRE Shadow Removal Challenge. MeanFlow represents the latest advance in flow-based methods, achieving state-of-the-art (SOTA) performance in one-step image generation. Its flexible diffusion process enables its direct application to image restoration. All results were either quoted from original publications or reproduced using their official implementations, with Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) as the primary evaluation metrics.